# Identifying Super Spreaders in Multilayer Networks

### v1

@@ 1 Introduction
Spreading phenomena in networks are investigated across various domains, including viral marketing, transportation systems, epidemiology, or social media [37]. These studies typically assume the propagation of a certain phenomenon over a graph-like structure simulating a real-world system. A key challenge in this context is to identify the agents that have the greatest impact on the spread. The significance of this task is underscored by its diverse applications, such as detecting vulnerable nodes in power networks, scheduling targeted immunisations, or selecting influencers for advertising campaigns. Nonetheless, addressing such problems requires accurate models of the systems under consideration. To this end, multilayer networks have been proposed [23]. These models provide a more nuanced representation of relationships by enabling, for instance, social ties to be split across multiple layers that correspond to different modes of interaction (e.g., physical, virtual). Considering spreading phenomena within multilayer setting can better capture the complexity of real-world systems, leading to more precise results.
@@ 1.1 Related works
Graph Neural Networks
Graph Neural Networks (GNNs) [46] represent a category of deep learning techniques specifically developed for inference on graph-structured data. The standard architecture of GNNs involves an iterative process of exchanging transformed node features and aggregating information from neighbouring nodes (so called message-passing). In a GNN with k layers, each node incorporates information from its k-hop neighbourhood. Variations in the design of the aggregation function and the combination function give rise to different GNN architectures [22, 7]. The resulting high-level node or graph representations are employed across a range of tasks. GNNs have been successfully applied to problems such as estimating information diffusion [5], graph-based reasoning [26], identifying the source of diffusion on graphs [44, 27] or even deep graph generation [28]. By leveraging proper information fusion [2], GNNs present a powerful solution to model graph-based data even on multilayer networks.
Influence Maximisation Problem
The problem of influence maximisation was originally introduced as a combinatorial optimisation task by [21], laying the foundation for a substantial body of research and practical applications over the following decades. It aims to identify a set of agents which, when used as diffusion seeds, maximise the spread of a given process across the network. The majority of traditional (i.e., non-learning-based) approaches can be classified into three categories: simulation-based, proxy-based, and heuristic-based. These methods often rely on centrality-based heuristics, utilising them directly, e.g. selecting top-$k$ nodes with the highest degree centrality, or building upon them, e.g. with discounting algorithms [6]. Although computationally efficient and interpretable, they operate under local assumptions that limit their adaptability to complex diffusion dynamics. For a comprehensive overview of these techniques, we refer the reader to a survey [39].
@@ 1.2 Assumptions and main contributions
In this study, we employ GNNs to model latent diffusion dynamics in multilayer networks and develop an end-to-end framework for identifying super-spreaders. This task can be viewed as a variant of the influence maximisation problem, which seeks to identify those agents within a graph that, if used as a single element seed sets, are able to disseminate information in the most effective manner — either by reaching the largest portion of the network or delivering it in the shortest possible time. Super-spreaders play a role analogous to that of a playmaker in a football team: if such a player is removed or underperforms, the team’s structure becomes destabilised and its winning potential is immediately diminished.
@@
Next, we propose a novel approach for identifying super-spreaders in multilayer networks by introducing the TopSpreadersNetwork (ts-net) — a robust GNN capable of generalising to previously unseen data, a feature rarely found among competitive trainable solutions. This ability stems from an architectural design that enables training on a diverse collection of networks, allowing the model to capture transferable structural patterns. We demonstrate the practical applicability and scalability of the proposed solution by predicting top-spreaders in real-world networks of varying sizes, with no constraints on the number of actors or network layers. Furthermore, a comparison with both classical heuristics and state-of-the-art trainable models shows that ts-net consistently outperforms competitive solutions. Source code developed for this study is available at: https://github.com/network-science-lab
@@ 2.1 Multilayer Networks
In essence, a multilayer network comprises a collection of intraconnected single-layer networks, where each actor may be represented by at most one node per layer, i.e. $V\subseteq A\times L$. For example, the representation of actor $a_{1}$ in layer $l_{2}$ is given by the node $v_{1}^{2}$. The model disallows interlayer edges, meaning that connections can occur only between nodes within the same layer: $(v_{1}^{1},v_{2}^{2})\notin E$. Furthermore, to simplify the analysis, all edges are assumed to be undirected and unweighted, that is, $(v_{1}^{1},v_{2}^{1})\equiv(v_{2}^{1},v_{1}^{1})$. At this point, it is worth noting that the concept of multilayer networks partially overlaps with that of heterogeneous graphs [38], the latter allowing a richer representation of the actor set (e.g. distinguishing between different classes of nodes) and permitting interlayer connections.
@@ 2.2 Top-spreaders identification
It is important to note that, by introducing ground-truth ranking $\mathbf{R}$ and basing subsequent parts of the work on this notion (especially the methods used to construct ground-truth rankings), we deviate from the classic influence maximisation problem. This departure is justified by the non-additivity of seed sets: the top-$k$ actors or nodes that are most effective at spreading information individually do not necessarily constitute the optimal $k$-element seed set [21].
Another important remark concerns the comparison between the ground-truth $\mathbf{R}$ and the predicted ranking $\mathbf{\hat{R}}$. A natural approach is to employ standard ranking-quality metrics such as Average Precision@$k$, or even the Jaccard index@$k$. However, in the problem we aim to address, the primary focus lies on the accuracy of the topmost predictions, where the most influential spreaders are expected to appear. Consequently, such metrics may be overly strict in evaluating discrepancies at the top of the ranking. For example, if the predicted ranking places the third and fourth most influential actors in the top two positions, it will yield the same Jaccard index as predicting two completely irrelevant actors in those positions, despite the former being a much more acceptable error.
To address this limitation and ensure a fairer comparison, we introduce the notion of a cumulated spreading score for the top-$k$ spreaders, as formalised in Def. 4:
@@
Based on the above, we can assess the quality of a predicted ranking $\mathbf{\hat{R}}$ in relation to the ground truth $\mathbf{R}$ (Def. 5). Instead of evaluating the prediction solely by the ordering of $\mathbf{\hat{R}}$, we compare the spreading scores of the top-$k$ actors in $\mathbf{\hat{R}}$ with those of the top-$k$ actors in $\mathbf{R}$.
@@
The relative cumulated score of a predicted top-$k$ actors from a ranking $\mathbf{\hat{R}}$ w.r.t. the ground-truth ranking $\mathbf{R}$ is given by:
@@
The task of identifying the top-$k$ spreaders in a network $M$, given a ground-truth ranking $R$, can be formulated as finding a function $\phi$ satisfying the following criterion:
@@ 3.2 Simulating information spreading
The second step involved simulating diffusion under the Independent Cascade Model (ICM) [14], adapted to account for the multilayer structure of the graph (MICM). Each actor in the network was used as a single-element seed set. During each simulation, we recorded four metrics that capture the spreading potential of the initiating actor, following the approach in [41]. These form a spreading-potential vector, as defined in Def. 7.
@@
In total, we performed simulations under forty distinct configurations of MICM, with each experiment on particular diffusion setup repeated forty times to reduce the stochastic effects inherent in the propagation model. Configurations of the MICM were defined by the Cartesian product of the two protocol functions: $\delta\in\{AND,OR\}$, and twenty activation probabilities: $\pi\in\{0.05,0.10,\ldots,1.00\}$. It is worth noting that we adopted the cascade-based model, rather than LTM, as it allows diffusion to be initiated from a single actor — an essential requirement for computing individual spreading potentials.
@@ 3.3 Defining spreading-potential score
A score function that maps vectors of spreading potentials to a value in range $[0,1]$ is defined as:
@@
The rationale behind this formulation reflects our interpretation of what constitutes a super-spreader. We prioritise the number of exposed actors, assigning it the highest weight, as it is typically regarded as the most critical metric in works related to spreading phenomena. All the remaining metrics are weighted equally, as none of them is deemed dominant. Moreover, our intention is to reward actors who spread information rapidly. Hence, we minimise $\tilde{p}_{sl}$ and $\tilde{p}_{pl}$ to favour fast and early spreaders.
@@ 3.4 Final metrics used in the study
This analysis led to the selection of the final metrics used to evaluate the performance of our proposed model and baseline methods. Primarily, we assessed the relative $sps$ of the actor identified as the most influential spreader in the network. Additionally, to account for the stratified nature of the $sps$ distribution, we defined a saddle ground-truth point at the 80th percentile for each network, enabling the evaluation of prediction quality for top spreaders. To assess the robustness of the methods, we also considered the area under the cumulative relative score curve, both up to the saddle point and over the entire range. For clarity, these metrics are summarised in Tab. 2. Finally, we note a technical detail: to ensure consistency, all distributions were normalised by the number of actors, resulting in relative prediction curves confined to the $[0,1]$ range on both axes.
Table 2: Final metrics, bounded within the range $[0,1]$, used to assess the effectiveness of methods for identifying top spreaders.
@@
$S_{auc}$ $\int_{1/|A|}^{s}y_{rel}(\mathbf{R},\mathbf{\hat{R}},k)$ Area under the cumulative relative score curve up to the saddle point
$S_{val}$ $y_{rel}(\mathbf{R},\mathbf{\hat{R}},s)$ Cumulative relative score of the identified top spreaders at the saddle point
$F_{auc}$ $\int_{1/|A|}^{|A|}y_{rel}(\mathbf{R},\mathbf{\hat{R}},k)$ Area under the full cumulative relative score curve
@@ 4 Proposed model – TopSpreadersNetwork
One of the central contributions of this study is a supervised, inductive graph neural network model: ts-net. The model comprises approximately $458{,}000$ trainable parameters and is designed to predict a vector of spreading potentials $\mathbf{\hat{p}}$, which is subsequently used to identify super-spreaders in size-adaptive, multilayer networks. The core concept of the model is illustrated in Fig. 2. Owing to the adopted design choices, ts-net produces robust predictions across networks of varying size, density, and structural complexity.
@@ 4.1 Structural design
The encoder is shared across all layers of the processed network and is responsible for computing actor embeddings independently for each layer. It processes the layer-specific adjacency matrix $\mathbf{A}$, the and a feature vector $\mathbf{x}$. This module consists of a four-layer hybrid GNN architecture that combines GIN [48] and GAT [43] blocks, enhanced with dropout [17], batch normalisation [18], and LeakyReLU [31] activation. The hybrid design enables the model to capture rich structural context while maintaining strong generalisation capabilities, particularly in networks with complex topologies aligning with the intuition presented in [34] and supported by the experimental results (see App. A).
@@ 4.2 Techniques for robust prediction
This section outlines the key design choices and techniques incorporated to enhance the predictive capabilities of the ts-net. An overview of the data processing flow during both training and inference is provided in Fig. 3. While the training procedure largely follows standard supervised learning paradigms, several non-trivial adaptations were introduced to improve model performance and generalisation. These are detailed in the following subsections.
@@
Since the employed diffusion mechanism considers solely actor’s connections across the network, no other actor attributes are meaningful in the context of this problem. Therefore, we use zero vectors as actor features, i.e. $\mathbf{x}=\mathbf{0}\in\mathbb{R}^{|A|\times d_{in}}$. As shown in App. A, this choice is supported by experiments involving alternative feature sets. Moreover, motivated by insights from the dataset analysis (Sec. 3.4), we introduced a data transformation module into the training pipeline to enhance the visibility of super-spreaders. Specifically, we applied a function $transform:\mathbf{p}\rightarrow\mathbf{p}_{transf}$ to the ground-truth labels used during training. After evaluating the effectiveness of various transformation functions (see App. A), we selected one referred to as $scatter$, which amplifies the scores of high-potential spreaders while attenuating those of weaker ones. The transformation is defined as $\mathbf{p}_{transf}=\exp(3\cdot\mathbf{p})/\exp(3)$. Importantly, this function preserves interpretability, as it is invertible and allows for reconstruction of the original target values.
@@ 5 Evaluation and results
A comprehensive set of experiments was designed to evaluate the effectiveness of ts-net with respect to the following questions: (i) how does the proposed method perform on both synthetic and real-world multilayer networks compared to competitive solutions (ii) can an inductive approach achieve performance comparable to that of state-of-the-art approaches (iii) what are the limitations of the examined methods in handling structurally diverse graphs (iv) how do the individual components of ts-net contribute to its overall performance (v) how do different hyperparameter settings, tasks, and dataset variants affect the performance of ts-net? In the following section, we provide answers to most of these questions, while referring interested readers to App. A for a detailed discussion of the two latter ones.
@@ 5.1 Experimental setup
To provide a comprehensive comparison of the capabilities of ts-net, we selected seven competitive solutions, encompassing both heuristic methods and learning-based models. It is important to note that all selected methods address the broader problem of influence maximisation, rather than the more specific task of identifying top spreaders. Moreover, some were originally developed for homogeneous networks and thus required adaptation to the multilayer context. Our selection was intended to reflect a broad spectrum of successful, state-of-the-art approaches that are most closely related in objective and scope to the problem under consideration.
The heuristic group includes classic centrality measures: Degree Centrality (deg-c) and One-Hop Neighbourhood Size [32] (nghb-s), along with their discounting counterparts: Degree Centrality Discount [6] (deg-cd) and Neighbourhood Size Discount [9] (nghb-sd). These were chosen for their low computational cost and ability to yield actor rankings compatible with multilayer settings. We also include a random baseline for reference. For learning-based models, we selected DeepIM [29] due to its data-driven design and recent impact in influence maximisation on homogeneous graphs. Despite its transductive nature, it serves as a strong neural baseline. As a representative of unsupervised embedding-based techniques adaptable to multilayer data we selected Multi-node2vec [45] with KMeans clustering (mn2v-km) to represent unsupervised embedding-based techniques adaptable to multilayer data.
The experimental procedure consists of three stages. First, ts-net, being the only inductive method among those evaluated, is trained on synthetic graphs, with $168$ used for training and $20$ for validation. In the second stage, all methods are applied to generate rankings of super spreaders on both synthetic and real-world networks not seen by the ts-net. For transductive methods, training is performed individually on each test network at this stage. Finally, the resulting rankings are evaluated using the metrics listed in Tab. 2. We refer the reader to App. B for results based on standard ranking metrics, which, however, are not substantially different from those presented below.

### v2

@@ 1 Introduction
Spreading phenomena in networks are investigated across various domains, including viral marketing, transportation systems, epidemiology, or social media [38]. These studies typically assume the propagation of a certain phenomenon over a graph-like structure simulating a real-world system. A key challenge in this context is to identify the agents that have the greatest impact on the spread. The significance of this task is underscored by its diverse applications, such as detecting vulnerable nodes in power networks, scheduling targeted immunisations, or selecting influencers for advertising campaigns. Nonetheless, addressing such problems requires accurate models of the systems under consideration. To this end, multilayer networks have been proposed [24]. These models provide a more nuanced representation of relationships by enabling, for instance, social ties to be split across multiple layers that correspond to different modes of interaction (e.g., physical, virtual). Considering spreading phenomena within a multilayer setting can better capture the complexity of real-world systems, leading to more precise results.
@@ 1.1 Related works
Graph neural networks.
Graph Neural Networks (GNNs) [47] represent a category of deep learning techniques specifically developed for inference on graph-structured data. The standard architecture of GNNs involves an iterative process of exchanging transformed node features and aggregating information from neighbouring nodes (so called message-passing). In a GNN with k layers, each node incorporates information from its k-hop neighbourhood. Variations in the design of the aggregation function and the combination function give rise to different GNN architectures [23, 7]. The resulting high-level node or graph representations are employed across a range of tasks. GNNs have been successfully applied to problems such as estimating information diffusion [5], graph-based reasoning [27], identifying the source of diffusion on graphs [45, 28], or even deep graph generation [29]. By leveraging proper information fusion [2], GNNs present a powerful solution to model graph-based data even on multilayer networks.
Influence maximisation problem.
[…]
@@ 1.2 Assumptions and main contributions
In this study, we employ GNNs to model latent diffusion dynamics in multilayer networks and develop an end-to-end framework for identifying super-spreaders. This task can be viewed as a variant of the influence maximisation problem, which seeks to identify those agents within a graph that, if used as a single element seed set, are able to disseminate information in the most effective manner — either by reaching the largest portion of the network or delivering it in the shortest possible time. Super-spreaders play a role analogous to that of a playmaker in a football team: if such a player is removed or underperforms, the team’s structure becomes destabilised and its winning potential is immediately diminished.
@@
Next, we propose a novel approach for identifying super-spreaders in multilayer networks by introducing the TopSpreadersNetwork (ts-net) — a robust GNN capable of generalising to previously unseen data, a feature rarely found among competitive trainable solutions. This ability stems from an architectural design that enables training on a diverse collection of networks, allowing the model to capture transferable structural patterns. We demonstrate the practical applicability and scalability of the proposed solution by predicting top-spreaders in real-world networks of varying sizes, with no constraints on the number of actors or layers. Furthermore, a comparison with both classical heuristics and state-of-the-art trainable models shows that ts-net consistently outperforms competitive solutions. The source code is available at: https://github.com/network-science-lab/infmax-trainer-icm-mln, appendices are published as [10].
@@ 2.1 Multilayer networks
In essence, a multilayer network comprises a collection of intraconnected single-layer networks, where each actor may be represented by at most one node per layer, i.e., $V\subseteq A\times L$. For example, the representation of actor $a_{1}$ in layer $l_{2}$ is given by the node $v_{1}^{2}$. The model disallows interlayer edges, meaning that connections can occur only between nodes within the same layer: $(v_{1}^{1},v_{2}^{2})\notin E$. Furthermore, to simplify the analysis, all edges are assumed to be undirected and unweighted, that is, $(v_{1}^{1},v_{2}^{1})\equiv(v_{2}^{1},v_{1}^{1})$. At this point, it is worth noting that the concept of multilayer networks partially overlaps with that of heterogeneous graphs [39], the latter allowing a richer representation of the actor set (e.g., distinguishing between different classes of nodes) and permitting interlayer connections.
@@ 2.2 Top-spreaders identification
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.2 Simulating information spreading
The second step involved simulating diffusion under the Independent Cascade Model (ICM) [15], adapted to account for the multilayer structure (MICM). Each actor in the network was used as a single-element seed set. During each simulation, we recorded four metrics that capture the spreading potential of the initiating actor, following [42]. These form a spreading-potential vector (Def. 7).
@@
In total, we performed simulations under forty distinct configurations of MICM, with each experiment on a particular diffusion setup repeated forty times to reduce the stochastic effects inherent in the propagation model. Configurations of the MICM were defined by the Cartesian product of the two protocol functions: $\delta\in\{AND,OR\}$, and twenty activation probabilities: $\pi\in\{0.05,0.10,\ldots,1.00\}$. It is worth noting that we adopted the cascade-based model, rather than LTM, as it allows diffusion to be initiated from a single actor — an essential requirement for computing individual spreading potentials.
@@ 3.3 Defining spreading-potential score
A score function that maps vectors of spreading potentials to a value in the range [0,1][0,1][ 0 , 1 ] is defined as:
@@
The rationale behind this formulation reflects our interpretation of what constitutes a super-spreader. We prioritise the number of exposed actors, as it is typically regarded as the most critical metric in works related to spreading phenomena. All the remaining metrics are weighted equally, as none of them is deemed dominant. Moreover, our intention is to reward actors who spread information rapidly. Hence, we minimise $\tilde{p}_{sl}$ and $\tilde{p}_{pl}$ to favour fast and early spreaders. These weights, however, could also be arranged differently depending on the task. For instance, if one is concerned with abrupt diffusion, greater weight may be assigned to the activation peak ($\tilde{p_{pi}}$).
@@ 3.4 Final metrics used in the study
This analysis led to the selection of the final metrics used to evaluate the performance of our proposed model and baseline methods. Primarily, we assessed the relative $sps$ of the actor identified as the most influential spreader in the network. Additionally, to account for the stratified nature of the $sps$ distribution, we defined a saddle ground-truth point $s$ at the 80th percentile for each network, enabling the evaluation of prediction quality for top spreaders. To assess the robustness of the methods, we also considered the area under the cumulative relative score curve, both up to the saddle point and over the entire range. For clarity, these metrics are summarised in Tab. 2. Finally, we note a technical detail: to ensure consistency, all distributions were normalised by the number of actors (i.e., $k\in\{1,\dots,|A|\}$ was transformed into $\tilde{k}\in\{\frac{1}{|A|},\dots,1\}$), resulting in relative prediction curves confined to the [0,1][0,1][ 0 , 1 ] range on both axes.
Table 2: Final metrics, bounded within the range [0,1][0,1][ 0 , 1 ], used to assess the effectiveness of methods for identifying top spreaders.
@@
$S_{auc}$ $\sum_{\frac{1}{|A|}\leq\tilde{k}\leq s}y_{rel}(\mathbf{R},\mathbf{\hat{R}},\tilde{k})$ Area under the relative cumulated score curve up to the saddle point
$S_{val}$ $y_{rel}(\mathbf{R},\mathbf{\hat{R}},s)$ Relative cumulated score of the identified top spreaders at the saddle point
$F_{auc}$ $\sum_{\frac{1}{|A|}\leq\tilde{k}\leq 1}\hskip-8.00003pty_{rel}(\mathbf{R},\mathbf{\hat{R}},\tilde{k})$ Area under the full relative cumulated score curve
@@ 4 Proposed model – TopSpreadersNetwork
One of the central contributions of this study is a supervised, inductive graph neural network model: ts-net. The model comprises approximately 458,000458{,}000458 , 000 trainable parameters and is designed to predict a vector of spreading potentials $\mathbf{\hat{p}}$, which is subsequently used to identify super-spreaders in size-adaptive, multilayer networks. The core concept of the model is illustrated in Fig. 2. Owing to the adopted design choices, ts-net produces robust predictions across networks of varying size, density, and structural complexity.
@@ 4.1 Structural design
The encoder is shared across all layers $L$ of the processed network $M$ and is responsible for computing actor embeddings independently for each layer $l\in L$. It processes the layer-specific adjacency matrix $\mathbf{A}$ and a feature vector $\mathbf{x}$. This module consists of a four-layer hybrid GNN architecture that combines GIN [49] and GAT [44] blocks, enhanced with dropout [18], batch normalisation [19], and LeakyReLU [32] activation. The hybrid design enables the model to capture rich structural context while maintaining strong generalisation capabilities, particularly in networks with complex topologies, aligning with the intuition presented in [35] and supported by the experimental results (see App. A).
@@ 4.2 Techniques for robust prediction
This section outlines the key design choices made to improve ts-net’s predictive capabilities. Fig. 3 presents the data processing flow during training and inference. Although the former largely follows standard supervised learning, several non-trivial adaptations were introduced to improve model performance and generalisation.
@@
Since the employed diffusion mechanism considers solely actor’s connections across the network, no other actor attributes are meaningful in the context of this problem. Therefore, we use zero vectors as actor features, i.e., $\mathbf{x}=\mathbf{0}\in\mathbb{R}^{|A|\times d_{in}}$. As shown in App. A, this choice is supported by experiments involving alternative feature sets. Moreover, motivated by insights from the dataset analysis (Sec. 3.4), we introduced a data transformation module into the training pipeline to enhance the visibility of super-spreaders. Specifically, we applied a function $transform:\mathbf{p}\rightarrow\mathbf{p}_{transf}$ to the ground-truth labels used during training. After evaluating the effectiveness of various approaches (see App. A), we selected one referred to as $scatter$, which amplifies the scores of high-potential spreaders while attenuating those of weaker ones. The transformation is defined as $\mathbf{p}_{transf}=\exp(3\cdot\mathbf{p})/\exp(3)$. Importantly, this function preserves interpretability, as it is invertible and allows for reconstruction of the original target values.
@@ 5 Evaluation and results
A comprehensive set of experiments was designed to evaluate the effectiveness of ts-net with respect to the following questions: (i) how does the proposed method perform on both synthetic and real-world multilayer networks compared to competitive solutions (ii) can an inductive approach achieve performance comparable to that of state-of-the-art approaches (iii) what are the limitations of the examined methods in handling structurally diverse graphs (iv) how do the individual components of ts-net contribute to its overall performance (v) how do different hyperparameter settings, tasks, and dataset variants affect the performance of ts-net? In the following section, we provide answers to most of these questions, while referring interested readers to App. A for a detailed discussion of the latter two.
@@ 5.1 Experimental setup
To provide a comprehensive comparison of the capabilities of ts-net, we selected seven competitive solutions, encompassing both heuristic methods and learning-based models. Notably, all of them address the broader problem of influence maximisation, rather than the more specific task of identifying top spreaders. Moreover, some were originally developed for homogeneous networks and thus required adaptation to the multilayer context. Our choice was intended to reflect a broad spectrum of successful, state-of-the-art methods that are most closely related in objective and scope to the considered problem.
The heuristic group includes classic centrality measures: Degree Centrality (deg-c) and One-Hop Neighbourhood Size [33] (nghb-s), along with their discounting counterparts: Degree Centrality Discount [6] (deg-cd) and Neighbourhood Size Discount [9] (nghb-sd). These were chosen for their low computational cost and ability to yield actor rankings compatible with multilayer settings. We also include a random baseline for reference. For learning-based models, we selected DeepIM [30] due to its data-driven design and recent impact in influence maximisation on homogeneous graphs. Despite its transductive nature, it serves as a strong neural baseline. As a representative of unsupervised embedding-based techniques adaptable to multilayer data, we selected Multi-node2vec [46] with KMeans clustering (mn2v-km) to represent unsupervised embedding-based techniques adaptable to multilayer data.
The experimental procedure consists of three stages. First, ts-net, being the only inductive method among those evaluated, is trained on synthetic graphs, with 168168168 used for training and 202020 for validation. In the second stage, all methods are applied to generate rankings of super spreaders on both synthetic and real-world networks not seen by the ts-net. For transductive methods, training is performed individually on each test network at this stage. Finally, the resulting rankings are evaluated using the metrics listed in Tab. 2. We refer the reader to App. B for results based on standard ranking metrics, which, however, are not substantially different from those presented below.